OpenStudy (anonymous):

solve the system of equations 3x+2y=2 -2x+y=8

OpenStudy (micahm):

freemathhelp.com

OpenStudy (anonymous):

3x+2y = 2 -2x + y = 8 y=8+2x Put y in first equation

OpenStudy (blurbendy):

If you look at the bottom equation, you can solve y pretty easily y = 8 + 2x Now, substitute 8 + 2x in for y in the first equation. what do you get?